# Graph Neural Networks for Jamming Source Localization

### v1

@@ 5.1 Static Evaluation for Jamming Localization
To analyze the impact of node arrangements and sampling coverage on localization accuracy, we evaluate performance in a static setting where devices remain fixed at predefined locations. Each instance consists of nodes randomly placed within the geographic area $\mathcal{A}=\{(x,y)\in\mathbb{R}^{2}\mid 0\leq x,y\leq 1500\}$, following circular, triangular, rectangular, or uniformly random layouts. These configurations define the sampling region $\mathcal{R}\subseteq\mathcal{A}$, where noise floor measurements are collected. The jammer’s position $\mathbf{x}_{j}$ is also randomly assigned within $\mathcal{A}$, independently of node placement, resulting in two distinct localization scenarios: when $\mathbf{x}_{j}\in\mathcal{R}$, proximity to source yields more precise sampling of interference, whereas when $\mathbf{x}_{j}\in\mathcal{A}\setminus\mathcal{R}$, localization relies on extrapolation from peripheral observations.
@@
Table 1 presents the jammer localization performance across various methods in the static experiment. The classical localization methods exhibit significantly higher errors across all sampling scenarios. Among them, WCL performs best with an overall RMSE of 143.1m, while LSQ yields the worst performance. The suboptimal performance of path-loss-based approaches (PL, MLE, MLAT, LSQ) is likely attributed to their dependency on estimating the path loss exponent ($\gamma$) and jammer transmit power ($P_{t}^{\text{jam}}$) in order to estimate the jammer position [12]. The GNN-based approaches consistently outperform these classical methods. CAGE delivers the highest overall performance, achieving an RMSE of 66.6m, followed by GAT with an RMSE of 74.4m, improving localization accuracy as compared to WCL by 114.9%.
@@
A significant performance gap occurs depending on whether the jammer is located inside or outside the sampled region $\mathcal{R}$. Localization outside $\mathcal{R}$ is inherently more challenging due to extrapolation beyond measured signals. Classical methods like WCL perform well within densely and radially symmetric sampled regions [25], resulting in an RMSE of 53.3m (inside). However, their performance sharply declines when extrapolating, with RMSE increasing to 232.8m (outside). In contrast, GAT demonstrates better robustness, with RMSE rising moderately from 40.0m (inside) to 108.8m (outside). CAGE achieves the lowest errors overall, with RMSE of 35.7m (inside) and 97.5m (outside). Notably, the circular topology significantly enhances performance when the jammer is outside $\mathcal{R}$ by maximizing angular diversity and preserving directional information. These findings highlight the critical role of spatial coverage and the superior robustness of GNN-based methods for long-range jammer localization.
@@
Table 2: Jammer localization results for dynamic trajectories, reporting MAE and RMSE over three trials. Results are categorized by distance intervals based on the minimum distance $d$ (m) between the jammer and the nearest sampled measurement. The final column reports the overall mean across the entire trajectory, computed considering each $s_{i}^{t}$ without skipping intermediate measurements.
Method Distance to the Jammer Mean
@@ 5.2 Dynamic Evaluation for Jamming Localization
We evaluate localization performance in a dynamic 3D environment where a mobile device navigates through space while tracking a jamming source. The trajectory follows an encirclement pattern, where the device gradually converges toward the jammer while collecting measurements from varying distances and angles. This experiment enables the analysis of localization performance under varying observational constraints, evaluating how well methods estimate the jammer’s position with limited initial data and how accuracy evolves as additional spatially diverse measurements are incorporated. During training, we employ random cropping of trajectory segments to expose the model to varied subsequences of the jamming encounter, enhancing robustness to partial observations and improving generalization across different trajectory lengths.
@@
Figure 3: Localization accuracy along a single trajectory for CAGE. The red cross-mark represents the jammer position, the orange mark represents the WCL reference prediction, and the blue mark represents the model prediction.
Table 2 presents the jammer localization performance across various methods in the dynamic experiment. The results are categorized by distance intervals, where each column represents a different proximity range between the jammer and the nearest sampled measurement. The final column reports the overall mean across trajectories. Notably, WCL achieves an RMSE of 11.6m in the closest range ($d\in[50,0]$). However, at larger distances ($d>500$), WCL degrades significantly with an RMSE of 372.4m. CAGE leverages confidence-weighted estimation to adaptively integrate WCL’s structured prior with the GNN-based predictions, resulting in the most consistent and robust performance. With an RMSE of 5.7m in $d\in[50,0]$, CAGE significantly outperforms both WCL and GAT. Interestingly, while CAGE integrates WCL as a prior, its benefits extend beyond dense, symmetric scenarios. The performance gap is particularly evident in sparse sampling conditions, where CAGE maintains significantly lower error than WCL and GAT, highlighting its ability to generalize beyond the expected advantages of structured priors.
Figure 2 shows localization performance trends over varying distances to the jammer and noise floor conditions. In Figure 2(a), WCL performs well at short distances, surpassing GAT only very close to the jammer ($\sim$20 m). GAT exhibits greater robustness over most of the trajectory, while CAGE consistently outperforms both methods. Figure 3 visually compares localization predictions along a trajectory, demonstrating progressive refinement after initial attack detection.
@@
5.3 Ablation Study on CAGE Components
Table 3 presents the ablation study on CAGE components. The columns $\mathcal{L}_{\text{GNN}}$ and $\mathcal{L}_{\text{Adapt}}$ indicate the use of graph-based and adaptive losses, respectively. “SN” and “SN Edges” denote the presence of a supernode and the edge type. “Con.” specifies the confidence layer type (linear or 3-layer MLP with ReLU and final sigmoid), while “Con. In” and “Reg. In” indicate the input sources for the confidence layer and regressor. “Con. Out” specifies whether a single or multiple confidence outputs are used. Finally, $\hat{\mathbf{x}}_{\text{final}}$ and $\hat{\mathbf{x}}_{\text{GNN}}$ represent the final estimated position and direct GNN prediction, respectively. The symbols , , and denote pooled graph representations with a supernode, without a supernode, and exclusively based on the supernode.
6 Conclusion
This study presents CAGE, a graph-based jammer localization framework that integrates attention-based GNNs with confidence-weighted adaptive estimation. By leveraging spatial priors and learned features, CAGE balances weighted centroid localization with graph-based learning for robust performance across diverse conditions. Evaluations and ablation studies show it outperforms classical and learning-based methods, especially in generalization beyond observed data and handling noisy and obscure signals.
Appendix 0.A Ablation Studies
@@
0.A.1 Node Features, Graph Construction, and Augmentations
Table 5 evaluates the impact of various node features in static and dynamic settings using the GCN model on a validation set. Incorporating Cartesian coordinates and weighted centroid distances improves RMSE, while azimuth-based features do not provide benefits. Local noise statistics, particularly mean noise deviation, significantly enhance accuracy. Furthermore, we incorporate unique IDs [18], with which we observe a degradation in performance, possibly due to a loss in the generalization ability of the model [32]. In dynamic scenarios, direction vectors and temporal noise variations further reduce RMSE, highlighting the importance of motion-aware attributes. Table 4 analyzes key graph construction and pooling elements. We find that a lower neighborhood size ($k=3$) yields best results, with increasing $k$ leading to oversmoothing [9]. Among global pooling strategies, max pooling achieves the best RMSE. For augmentations, DropNode [5] with a 0.2 drop rate leads to the best improvement. For more information on the augmentation experiments, refer to Appendix 0.A.
@@ 0.A.2 Data Augmentations
We also evaluate combined augmentation strategies where Crop+DN refers to cropping followed by DropNode, DN+Feat Noise represents DropNode followed by feature corruption, and Crop+Feat Noise denotes cropping followed by feature corruption. Table 7 illustrates that combining these augmentations does not significantly improve performance. The best result is achieved with DN + Feat Noise, reinforcing our decision to use DropNode with $p=0.2$ exclusively.
Table 7: Performance of combined augmentation strategies.
Aug Crop+DN DN+Feat Noise Crop+Feat Noise
RMSE 55.5±0.1 53.0±0.1 55.3±0.2
Appendix 0.B Analysis of Confidence Weighting in CAGE
We analyze how the $\mathcal{L}_{\text{CAGE}}$ loss shapes confidence weight assignment in CAGE. To do so, we compare two training setups: one using only $\mathcal{L}_{\text{Adapt}}$ and another incorporating the full $\mathcal{L}_{\text{CAGE}}$ loss. Confidence weights are plotted against the jammer’s distance to analyze how the choice of loss function affects the model’s reliance on the WCL prior versus GNN-based predictions. This experiment is conducted using dynamic experiment data, described in Section 5.2, where the trajectory follows an encirclement pattern, gradually converging toward the jammer while collecting measurements from varying distances and angles.
Refer to caption Refer to caption
Figure 4: Effect of training loss on confidence weighting in CAGE. (a) Training with $\mathcal{L}_{\text{Adapt}}$ causes over-reliance on WCL due to low GNN confidence. (b) Training with $\mathcal{L}_{\text{CAGE}}$ increases initial GNN confidence and sharpens the transition near the jammer.
When training solely with $\mathcal{L}_{\text{Adapt}}$ (Figure 4a), the confidence weights remain relatively low even at large distances, indicating that the model strongly relies on WCL across all regions, even where GNN-based predictions should be more informative. Since the final prediction is computed as a weighted combination in Equation (12), higher confidence weights ($\boldsymbol{\alpha}$) indicate greater reliance on the GNN-based prediction, while lower weights shift reliance toward WCL. Because $\mathcal{L}_{\text{Adapt}}$ only minimizes blended prediction error, it fails to enforce a direct learning signal for the GNN, causing it to act as a correction to WCL rather than as an independent predictor.
In contrast, when $\mathcal{L}_{\text{CAGE}}$ is used (Figure 4b), the confidence weights start much higher ($\alpha\approx 1$), showing that the model correctly prioritizes GNN-based predictions in sparse sampling scenarios. As the distance to the jammer decreases, confidence in WCL increases, suggesting that the model has learned to recognize when WCL becomes reliable due to dense sampling and radial symmetry at close distances, which is a characteristic of the data used for this experiment. This improvement stems from enforcing independent learning on $\mathcal{L}_{\text{GNN}}$ when predicting the jammer’s position. Ultimately, incorporating $\mathcal{L}_{\text{CAGE}}$ results in a more structured confidence-weighting mechanism, leading to improved robustness and localization accuracy.
Figure 4b also reveals that among all confidence weights, $\alpha_{2}$, which corresponds to $\sin(\theta)$ in the final estimation, remains slightly lower than the others, even at large distances. This indicates that the model retains a small but consistent reliance on the azimuth component of the WCL estimate, even when predominantly using the GNN-based prediction. The fact that $\alpha_{2}$ is consistently below the other confidence values suggests that while the model learns to prioritize the GNN overall, it still leverages the WCL azimuth as a stable directional reference. Fine-tuning $\lambda$ in Equation (14) allows for adjusting the model’s reliance on different components, with $\lambda=0$ used in this work.
Appendix 0.C Further Results and Analysis

### v2

@@ 5.1 Static Evaluation for Jamming Localization
To analyze the impact of node arrangements and coverage on localization accuracy, we evaluate performance in a static setting where devices remain fixed at predefined locations. Each instance consists of nodes randomly placed within the geographic area $\mathcal{A}=\{(x,y)\in\mathbb{R}^{2}\mid 0\leq x,y\leq 1500\}$, following circular, triangular, rectangular, or uniformly random layouts. These configurations define the sampling region $\mathcal{R}\subseteq\mathcal{A}$, where noise floor measurements are collected. The jammer’s position $\mathbf{x}_{j}$ is also randomly assigned within $\mathcal{A}$, independently of node placement, resulting in two distinct localization scenarios: when $\mathbf{x}_{j}\in\mathcal{R}$, proximity to the source yields more precise sampling of interference, whereas when $\mathbf{x}_{j}\in\mathcal{A}\setminus\mathcal{R}$, localization relies on extrapolation from peripheral observations.
@@
Table 1 presents the jammer localization performance across various methods in the static experiment as measured by RMSE. MAE results are provided in Appendix 0.B Table 8. The classical localization methods exhibit significantly higher errors across all sampling scenarios. Among them, WCL performs best with an overall RMSE of 143.1m, while LSQ yields the worst performance. The suboptimal performance of path-loss-based approaches (PL, MLE, MLAT, LSQ) is likely attributed to their dependency on estimating the path loss exponent ($\gamma$) and jammer transmit power ($P_{t}^{\text{jam}}$) in order to estimate the jammer position [13]. The GNN-based approaches consistently outperform these classical methods. CAGE delivers the highest overall performance, achieving an RMSE of 66.6m, followed by GAT with an RMSE of 74.4m, improving localization accuracy as compared to WCL by 114.9%.
@@
A significant performance gap is observed based on whether the jammer is located inside or outside the sampled region $\mathcal{R}$. Localization outside $\mathcal{R}$ is inherently more challenging due to extrapolation beyond measured signals. Classical methods like WCL perform well within densely and radially symmetric sampled regions [25], resulting in an RMSE of 53.3m (inside). However, the performance sharply declines when extrapolating, with RMSE increasing to 232.8m (outside). In contrast, GAT demonstrates better robustness, with RMSE rising moderately from 40.0m (inside) to 108.8m (outside). CAGE achieves the lowest errors overall, with RMSE of 35.7m (inside) and 97.5m (outside). Notably, the circular topology significantly enhances performance when the jammer is outside $\mathcal{R}$ by maximizing angular diversity and preserving directional information, as supported by the principles guiding the use of circular antenna arrays in direction-finding applications [7]. These findings highlight the critical role of spatial coverage and the superior robustness of GNN-based methods for long-range jammer localization.
@@
Table 2: RMSE in jammer localization along dynamic trajectories, reported across distance intervals to the jammer and averaged over three trials. The final column reports the mean RMSE across the full trajectory. MAE results are provided in Appendix 0.B Table 9.
Method Distance to the Jammer (m) Mean
@@ 5.2 Dynamic Evaluation for Jamming Localization
We evaluate localization performance in a dynamic 3D environment where a mobile device travels through space while tracking a jamming source. The trajectory follows an encirclement pattern, where the device converges toward the jammer while collecting measurements from varying distances and angles. This experiment enables the analysis of localization performance under varying observational constraints, evaluating how well methods estimate the jammer’s position with limited initial data and how accuracy evolves as additional spatially diverse measurements are incorporated. During training, we employ random cropping of trajectory segments to expose the model to varied subsequences of the jamming encounter, enhancing robustness to partial observations and improving generalization across different trajectory lengths.
@@
Figure 3: Localization accuracy along a single trajectory. The red cross-mark is the jammer position, orange is WCL, and blue is the CAGE prediction.
Table 9 presents the jammer localization performance for various methods in the dynamic experiment for MAE. RMSE results are provided in Appendix 0.B Table 2. The results are categorized by distance intervals, where each column represents a different proximity range between the jammer and the nearest sampled measurement. The final column reports the overall mean across trajectories. Notably, WCL achieves an RMSE of 11.6m in the closest range ($d\in[50,0]$). However, at larger distances ($d>500$), WCL degrades significantly with an RMSE of 372.4m. CAGE leverages confidence-weighted estimation to adaptively integrate WCL’s structured prior with the GNN-based predictions, resulting in the most consistent and robust performance. With an RMSE of 5.7m in $d\in[50,0]$, CAGE significantly outperforms both WCL and GAT. Interestingly, while CAGE integrates WCL as a prior, its benefits extend beyond dense, symmetric scenarios. The performance gap is particularly evident in sparse sampling conditions, where CAGE maintains significantly lower error than WCL and GAT, highlighting its ability to generalize beyond the expected advantages of structured priors.
Figure 2 shows localization performance trends over varying distances to the jammer and noise floor conditions. In Figure 2(a), WCL performs well at short distances, surpassing GAT only when very close to the jammer ($\sim$20 m). While GAT demonstrates greater robustness along most of the trajectory, its performance is consistently surpassed by CAGE. Figure 3 visually compares localization predictions along a trajectory, demonstrating progressive refinement after initial attack detection.
5.3 Ablation Study
5.3.1 CAGE Components
We perform an ablation study to assess the contribution of individual components in the CAGE architecture, as detailed in Table 3. The columns $\mathcal{L}_{\text{GNN}}$ and $\mathcal{L}_{\text{Adapt}}$ indicate the use of graph-based and adaptive losses, respectively, as defined in Section 4. “SN” and “SN Edges” denote the presence of a supernode and its edge connections. “Con.” represents the confidence weight ($\alpha$) layer (linear or 3-layer MLP with ReLU activations and final sigmoid), while “Con. In” and “Reg. In” indicate the input data for the confidence weight layer and regressor, respectively. “Con. Out” indicates whether a single shared confidence weight is used for all predicted coordinates or separate weights are produced for each. Finally, $\hat{\mathbf{x}}_{\text{final}}$ and $\hat{\mathbf{x}}_{\text{GNN}}$ represent the estimated position using the marked loss functions in the table and the GNN-only prediction, respectively. The symbols , , and denote pooled graph representations with the supernode, without the supernode, and exclusively using the supernode.
5.3.2 Confidence Weighting in CAGE
We analyze how the $\mathcal{L}_{\text{CAGE}}$ loss shapes confidence weight assignment in CAGE by comparing two training setups: one using only $\mathcal{L}_{\text{Adapt}}$ and another incorporating the full $\mathcal{L}_{\text{CAGE}}$ loss. Confidence weights are plotted against the jammer’s distance to assess how the choice of loss function influences the model’s reliance on the WCL prior versus GNN-based predictions. This experiment uses the dynamic trajectory data described in Section 5.2, where measurements are collected while encircling the jammer from varying distances and angles. This setup yields conditions where the GNN can best leverage wide spatial context at long ranges, while WCL becomes increasingly reliable at close distances due to the dense, radially symmetric sampling created by the encircling motion.
@@ 5.3.2 Confidence Weighting in CAGE
When trained solely with $\mathcal{L}_{\text{Adapt}}$ (Figure 4a), the model assigns relatively low confidence weights across all distances. This suggests a persistent reliance on WCL, even in regions where GNN-based predictions should dominate. Since $\mathcal{L}_{\text{Adapt}}$ optimizes the blended prediction without enforcing a direct learning signal for the GNN, the model tends to treat it as a correction mechanism rather than as an independent estimator. In contrast, training with $\mathcal{L}_{\text{CAGE}}$ (Figure 4b) yields high initial confidence in the GNN prediction ($\alpha\approx 1$), especially at large distances. As the distance to the jammer decreases, confidence in WCL increases, indicating that the model learns when the WCL prior becomes more reliable, particularly in the densely sampled, radially symmetric regions near the jammer. This improvement stems from the explicit GNN supervision provided by $\mathcal{L}_{\text{GNN}}$, resulting in a more structured confidence-weighting mechanism and improved localization robustness.
Refer to caption Refer to caption
Figure 4: Effect of training loss on confidence weighting in CAGE. (a) Training with $\mathcal{L}_{\text{Adapt}}$ causes over-reliance on WCL due to low GNN confidence. (b) Training with $\mathcal{L}_{\text{CAGE}}$ increases initial GNN confidence, sharpening near the jammer.
Figure 4b also reveals that among all confidence weights, $\alpha_{2}$, which corresponds to $\sin(\theta)$ in the final estimation, remains slightly lower than the others, even at large distances. This suggests that while the model generally prioritizes GNN-based predictions, it continues to rely on the WCL azimuth as a stable directional reference. Although we used $\lambda=0$ in this work, tuning $\lambda$ in Equation (14) may further improve performance by better balancing the model’s reliance on different components.
@@
(d) Performance of combined augmentation strategies.
Aug Crop+DN DN+Feat Noise Crop+Feat Noise RMSE 55.5±0.1 53.0±0.1 55.3±0.2
5.3.3 Graph Construction, Pooling, and Augmentations
We investigate how different graph construction strategies, pooling methods, and augmentation techniques influence model performance, as summarized in Table 4. A lower neighborhood size ($k=3$) yields the best RMSE, with larger $k$ leading to oversmoothing [10]. Among global pooling strategies, max pooling performs best. For augmentations, DropNode [5] at a 0.2 drop rate offers the greatest improvement, and the combination of DropNode and feature noise achieves the best performance overall. See Appendix 0.A for further results on feature engineering and augmentations.
5.3.4 Downsampling Strategies for High-Resolution Signal Graphs
High-frequency sampling in Software Defined Radio (SDR) systems produces dense graphs with thousands of nodes (approx. 6000 per instance in our dynamic dataset; see Appendix 0.D, Table 11). To reduce computational cost while preserving key signal characteristics, we apply downsampling prior to graph construction. We compare two methods: (1) window averaging, which operates on the raw sequence of samples by dividing it into $|V|$ segments based on sample count and averaging the position and noise values within each; and (2) spatial binning with noise filtering, which groups samples into fixed 1m3 spatial bins based on their positions, averages the position and noise values within each bin, and retains only the $|V|$ bins with the highest average noise, as high-noise regions are more informative [13].
Table 5: Ablation study of downsampling techniques on high-resolution dynamic path data under LDPL model. The retained number of nodes is denoted by $|V|$.
$|V|$ Window averaging Spatial binning with noise filtering
200 74.5 22.1 13.2 38.7 78.4 21.6 10.5 40.5
600 63.2 21.8 14.7 34.3 63.5 16.3 9.8 31.7
800 57.1 23.9 15.7 32.2 59.3 15.3 10.4 29.7
1000 54.6 23.4 16.5 31.4 54.5 13.7 9.8 27.4
Table 5 presents results from the downsampling ablation experiment. Each segment $t_{\ell,u}$ denotes a subinterval of the normalized trajectory, with $\ell$ and $u$ indicating the lower and upper bounds of the time fraction (e.g., $t_{0.8,1.0}$ corresponds to the final 20% of the path). The results indicate that spatial binning with noise filtering consistently outperforms window averaging, particularly in later segments. Following these results, we adopt the spatial boning with noise filtering with $|V|=1000$ for all our experiments.
6 Conclusion
This work presents CAGE, the first graph-based framework for jamming source localization that reformulates the problem as an inductive graph regression task. Our approach integrates attention-based GNNs with a confidence-weighted fusion mechanism that adaptively balances learned predictions with structured spatial priors. Experiments across diverse static and dynamic scenarios demonstrate that CAGE consistently outperforms both classical and learning-based baselines. Through detailed ablations, we show that our design choices significantly enhance model robustness and performance. Future work includes incorporating temporal GNNs, analyzing the impact of the confidence weighting parameter $\lambda$, and evaluating cross-domain generalization beyond simulated environments.
Appendix 0.A Ablation Studies
0.A.1 Node Features, Graph Construction, and Augmentations
Table 6 evaluates the impact of various node features in static and dynamic settings using the GCN model on a validation set. Incorporating Cartesian coordinates and weighted centroid distances improves RMSE, while azimuth-based features do not provide benefits. Local noise statistics, particularly mean noise deviation, significantly enhance accuracy. Furthermore, we incorporate unique IDs [19], with which we observe a degradation in performance, possibly due to a loss in the generalization ability of the model [32]. In dynamic scenarios, direction vectors and temporal noise variations further reduce RMSE, highlighting the importance of motion-aware attributes.
@@ 0.A.2 Data Augmentations
We also evaluate combined augmentation strategies where Crop+DN refers to cropping followed by DropNode, DN+Feat Noise represents DropNode followed by feature corruption, and Crop+Feat Noise denotes cropping followed by feature corruption. Table 4d illustrates that combining these augmentations does not significantly improve performance. The best result is achieved with DN + Feat Noise, reinforcing our decision to use DropNode with $p=0.2$ exclusively.
Table 8: MAE in jammer localization for static scenarios, averaged over three trials with different seeds. Results are split by sampling geometry.
C T R RD Mean C T R RD Mean
MAE WCL 37.7 40.2 27.2 35.7 35.2 138.6 169.0 170.1 164.2 160.5 97.9
PL 122.2 83.3 90.6 82.0 94.5 289.4 247.3 290.6 267.4 273.7 184.1
  MLE 70.5 65.5 63.4 67.4 66.7 208.4 197.9 203.5 230.3 210.0 138.4
MLAT 121.6 89.6 75.3 64.8 87.8 282.2 259.9 275.2 235.8 263.3 175.6
LSQ 237.9 138.5 92.0 94.1 140.6 410.5 329.7 351.1 310.9 350.6 245.6
MLP 37.1 31.3 24.7 29.5 30.7 60.1 78.2 78.5 84.1 75.2 53.0
GCN 35.5 30.7 25.9 35.3 31.9 57.5 75.5 77.4 85.4 74.0 53.0
  PNA 34.3 27.7 21.9 26.7 27.7 56.5 72.7 74.1 79.4 70.7 49.2
GAT 33.3 27.4 21.5 27.2 27.4 55.8 72.0 73.5 77.9 69.8 48.6
  CAGE 28.4 23.6 19.8 25.0 24.2 46.5 61.6 65.1 68.8 60.5 42.5
Table 9: MAE in jammer localization along dynamic trajectories, reported across distance intervals to the jammer and averaged over three trials. The final column reports the mean MAE across the full trajectory.
Method Distance to the Jammer Mean
$d>500$ $d\in[500,200]$ $d\in[200,100]$ $d\in[100,50]$ $d\in[50,0]$
MAE WCL 316.8 220.6 95.2 34.6 7.7 28.2
PL 323.8 275.3 153.5 74.0 28.9 56.3
MLE 245.1 130.0 54.2 36.5 1381.8 1006.7
MLAT 237.8 192.5 122.2 85.8 57.7 73.1
MLP 140.5±9.7 83.7±4.9 35.5±1.5 21.8±0.9 14.8±0.8 20.7±0.4
GCN 123.0±3.8 65.4±1.0 25.2±0.1 14.3±0.2 8.1±0.2 13.1±0.1
PNA 162.1±5.2 95.1±2.6 40.9±2.2 26.3±1.1 19.5±1.0 25.7±1.0
GAT 88.7±10.6 49.3±1.4 23.1±0.7 13.3±0.4 7.3±0.4 11.4±0.5
CAGE 73.1±0.7 36.7±0.9 17.1±0.6 9.9±0.1 4.2±0.1 7.6±0.1
Appendix 0.B Further Results and Analysis